# Chemical Diagnostics to Unveil Environments Enriched by First Stars

### v1

@@ 4.3 A unique absorber
The absorber J1018+0548 agrees with an enrichment dominated by PISN or other Pop III SNe in all the proposed diagnostics (Fig. 2-3). By applying a chi square diagnostics (see Appendix B), we looked for the models that best reproduce its abundance pattern (Fig. 4). The minimum $\chi^{2}_{\nu}=0.19$ corresponds to an enrichment driven by two Pop III SNe ($\rm E_{SN}=1.8\times 10^{51}$ erg). However, a low $\chi^{2}_{\nu}=$ 0.37, is also obtained when considering a pollution from a $\rm 253M_{\odot}$ PISN (contributing to 10% of the metals), and a Pop III SN with $\rm E_{SN}=1.2\times 10^{51}$ erg. Both models fit very well the abundance ratios of the five light elements measured in the J1018+0548 absorber. However they strongly differ for elements heavier than Ca, in particular [Zn/Fe] which is $\approx 1.5$ dex lower for massive PISN enrichment.
@@ 5 Conclusions
In this Letter, we present novel chemical diagnostics to isolate PISN-enriched environments from those imprinted by other Pop III SNe and normal Pop II stars. To this end, we exploited the simple and general parametric model by Vanni et al. (2023b). We show that environments enriched by a single PISN dwell in well defined zones of the proposed chemical diagnostics. Indeed they cover $>2$ dex in [X/Fe] but have characteristic values of [Al, Mg, Na/Si] and [Si, C/O], covering $<0.5$ dex. However, when the contribution of normal Pop II SNe is accounted for, regions predominantly imprinted by PISNe ($\geq 60\%$ of metals) grow and blend with those mainly enriched by other Pop III or Pop II SNe, creating the so-called confusion regions.
By comparing our model predictions with measurements in high-$z$ absorbers, we see that most systems in the literature have elemental abundances consistent with being mainly imprinted by normal Pop II SNe, in agreement with previous studies (e.g. Becker et al., 2012; Salvadori & Ferrara, 2012). Note that we do not account for dust depletion since most absorbers in our literature sample are metal-poor, meaning that dust should not affect substantially their chemical abundances (Vladilo et al., 2018; Saccardi et al., 2023). In Figs. 1, 2, 3, we show in which direction the predictions would shift if dust was present, using the prescriptions of Konstantopoulou et al. (2024).
Our analysis of the publicly available high-resolution X-Shooter spectra for two quasars presented by Christensen et al. (2023), reconcile the chemical abundances of two absorption systems with literature data, i.e., with an enrichment dominated by normal Pop II SNe. Only the system at $z=6.5625$ along the sightline to VDES J0020–3653, for which high resolution spectra are not available, remains consistent with a PISN pollution for the three measured elements. Conversely, the Galactic halo star observed at high-resolution by Xing et al. (2023, 12 elements) is in agreement with an exclusive PISN-enrichment in all considered diagnostics. Still, more chemical elements (e.g., C) need to be measured to confirm a pure pollution by a massive PISN (see also Koutsouridou et al., 2024).
@@ Appendix A Analysis of X-Shooter quasar spectra
[…]
The absorption lines have been fitted with Voigt profiles using the Python package Astrocook (Cupani et al., 2022). For the three analysed systems, we assumed that the Doppler parameters were the same for all observed transitions. In the case of the two systems along the spectrum of ULAS J1342+0928, we fixed the Doppler parameter to the minimum value allowed by the resolving power of the observations (see Sodini et al. subm.). Due to the limited NIR coverage of the X-Shooter spectrum with respect to the NIRSpec one, we do not cover the Mg $\scriptstyle\rm II$ doublet transitions for either system.
In the case of the system at $z=6.2897$ in the spectrum of UHS J0439+1628, Christensen et al. (2023) identify as $\scriptstyle\rm I$ 1302 Å̊$\mathring{\rm A}$ at $z=6.288$ an absorption feature which Davies et al. (2023) fit in the X-Shooter spectrum as a $\scriptstyle\rm IV$ doublet at $z=5.1299$. For the purpose of this paper and in order to compare with the results by Christensen et al. (2023), we derived a conservative upper limit on the $\scriptstyle\rm I$ column density fitting the observed absorption assuming the redshift and the Doppler parameter of the Si $\scriptstyle\rm II$ 1260 Å̊$\mathring{\rm A}$ absorption line of this system, i.e. assuming $z\approx 6.2897$. The very good quality of the X-Shooter spectrum allowed us also to measure the column density of a very weak Fe $\scriptstyle\rm II$ 2382 Å̊$\mathring{\rm A}$ absorption line and to determine an upper limit to the transition of Al $\scriptstyle\rm II$ 1670 Å̊$\mathring{\rm A}$. The absorption lines and the Voigt profile fit for the three studied systems are shown in Fig. 5, 6 and 7; while the measured column densities are reported in Table 2.
Appendix B Fitting procedure
$\chi^{2}=\sum_{i=1}^{N}\dfrac{(O_{i}-T_{i})^{2}}{\sigma^{2}_{O_{i}}+\sigma^{2}_{T_{i}}}+\sum_{i=N+1}^{N+U}\dfrac{(O_{i}-T_{i})^{2}}{\sigma^{2}_{T_{i}}}\times\Theta(O_{i}-T_{i})+\sum_{i=N+U+1}^{N+U+L}\dfrac{(O_{i}-T_{i})^{2}}{\sigma^{2}_{T_{i}}}\times\Theta(T_{i}-O_{i}),$ (B1)
[…]
To find the best model for the enrichment of the absorption system at $z=3.39$ in the sightline of the QSO J1018+0548 (Saccardi et al., 2023), we look for the minimum $\chi^{2}_{\nu}$ by considering all the models presented in Sec. 2 (see Fig. 1, 2, 3), i.e. by accounting for an enrichment: (i) uniquely driven by a single PISN or a Pop III SN with different mass/energy; (ii) mainly ($\geq 60\%$ metals) driven by a single PISN/other Pop III SN with the contribution of Pop II SNe; (iii) only partially ($\leq 50\%$ metals) driven by a single PISN/other Pop III SN with the contribution of Pop II SNe. Furthermore, we also consider an imprint from two Pop III SNe with various masses and energies, including PISNe. As shown in Fig. 4 these models provide the best fits.

### v2

@@ 4.3 A unique absorber
The absorber J1018+0548 agrees with an enrichment dominated by PISN or other Pop III SNe in all the proposed diagnostics (Fig. 2-3). By applying a chi square diagnostics (see Appendix C), we looked for the models that best reproduce its abundance pattern, i.e. with reduced chi square, $\chi^{2}_{\nu}$, lower than two times the minimum one (Fig. 4). The minimum $\chi^{2}_{\nu}=0.19$ corresponds to an enrichment driven by two Pop III SNe with equal explosion energy ($\rm E_{SN}=1.8\times 10^{51}$ erg) but very different masses ($\rm\approx 10$ and $100\rm M_{\odot}$, see Table 3). However, a low $\chi^{2}_{\nu}=$ 0.37, is also obtained when considering a pollution from a $\rm 253M_{\odot}$ PISN (contributing to 10% of the metals), and a $\rm\approx 30M_{\odot}$ Pop III SN with $\rm E_{SN}=1.2\times 10^{51}$ erg. Both models fit very well the abundance ratios of the five light elements measured in the J1018+0548 absorber. However they strongly differ for elements heavier than Ca, in particular [Zn/Fe] which is $\approx 1.5$ dex lower for massive PISN enrichment.
@@ 5 Conclusions
In this Letter, we present novel chemical diagnostics to isolate PISN-enriched environments from those imprinted by other Pop III SNe and normal Pop II stars. To this end, we exploited the simple and general parametric model by Vanni et al. (2023b). We show that environments enriched by a single PISN dwell in well defined zones of the proposed chemical diagnostics. Indeed they cover $>2$ dex in [X/Fe] but have characteristic values of [Al, Mg, Na/Si] and [Si, C/O], covering $<0.5$ dex. However, when the contribution of normal Pop II SNe is accounted for, regions predominantly imprinted by PISNe ($\geq 60\%$ of metals) grow and blend with those mainly enriched by other Pop III or Pop II SNe, creating the so-called confusion regions. Our predictions are very general and thus can be used to interpret the chemical properties of various environments across cosmic times: high-z emitting galaxies (e.g. D’Eugenio et al., 2023), distant gaseous absorbers in the sightline of QSOs or GRBs (see also Ma et al., 2017b) and ancient stellar fossils in our Local Group (Vanni et al., 2023b).
By comparing our model predictions with measurements in high-$z$ absorbers, we see that most systems in the literature have elemental abundances consistent with being mainly imprinted by normal Pop II SNe, in agreement with previous studies (e.g. Becker et al., 2012; Salvadori & Ferrara, 2012). Note that we do not account for dust depletion since most absorbers in our literature sample are metal-poor, meaning that dust should not affect substantially their chemical abundances (Vladilo et al., 2018; Saccardi et al., 2023). In Figs. 1, 2, 3, we show in which direction the predictions would shift if dust was present, using the prescriptions of Konstantopoulou et al. (2024) (see also De Cia et al., 2016).
Our analysis of the publicly available high-resolution X-Shooter spectra for two QSOs presented by Christensen et al. (2023), reconcile the chemical abundances of two absorption systems, suggesting exclusive PISN-enrichment, with literature data, i.e., with an enrichment dominated by normal Pop II SNe. Only the system at $z=6.5625$ along the sightline to VDES J0020–3653, for which high resolution spectra are not available, remains consistent with a PISN pollution for the three measured elements. Conversely, the Galactic halo star observed at high-resolution by Xing et al. (2023, 12 elements) is in agreement with an exclusive PISN-enrichment in all considered diagnostics. Still, more chemical elements (e.g., C) need to be measured to confirm a pure pollution by a massive PISN (see also Koutsouridou et al., 2024).
@@ Appendix A Analysis of X-Shooter quasar spectra
In this section, we report the results of the fit of three absorption systems discussed in Christensen et al. (2023) for which we analysed the corresponding X-Shooter spectra. The spectrum of UHS J0439+1634 is the one released by the XQR-30 collaboration (D’Odorico et al., 2023b), while the spectrum of ULAS J1342+0928 was reduced and analysed in this work. Briefly, ULAS J1342+0928 was observed with VLT/X-Shooter in 2017 and 2018 (PID 098.B-0537, 0100.A-0898) for $\sim 23$h adopting a slit of 0.6” in the NIR arm. We retrieved the raw frames from the ESO archive and reduced them with the custom pipeline described in D’Odorico et al. (2023b). The combined spectrum has a nominal resolving power $R\simeq 8100$ and a signal-to-noise ratio S/N $\simeq 19$ per 10 km s-1 bin at $\lambda=1285$̊$\mathring{\rm A}$ rest frame. This value, converted to the JWST/NIRSpec wavelength bin at $\lambda\sim 1.1$ $\mu$m ($\Delta v\simeq 64$ km s-1) increases to S/N $\simeq 50$, comparable to the one of the spectra analysed by Christensen et al. (2023, S/N $\approx 60$, see their Fig. 5).
The absorption lines have been fitted with Voigt profiles using the Python package Astrocook (Cupani et al., 2022). For the three analysed systems, we assumed that the Doppler parameters were the same for all observed transitions. In the case of the two systems along the spectrum of ULAS J1342+0928, we fixed the Doppler parameter to the minimum value allowed by the resolving power of the observations (see Sodini et al. subm.). However, we checked that decreasing the Doppler parameter to 6 km s-1 the results do not change significantly. Due to the limited NIR coverage of the X-Shooter spectrum with respect to the NIRSpec one, we do not cover the Mg $\scriptstyle\rm II$ doublet transitions for either system.
In the case of the system at $z=6.2897$ in the spectrum of UHS J0439+1634, Christensen et al. (2023) identify as $\scriptstyle\rm I$ 1302 Å̊$\mathring{\rm A}$ at $z=6.288$ an absorption feature which Davies et al. (2023) fit in the X-Shooter spectrum as a $\scriptstyle\rm IV$ doublet at $z=5.1299$. For the purpose of this paper and in order to compare with the results by Christensen et al. (2023), we derived a conservative upper limit on the $\scriptstyle\rm I$ column density fitting the observed absorption assuming the redshift and the Doppler parameter of the Si $\scriptstyle\rm II$ 1260 Å̊$\mathring{\rm A}$ absorption line of this system, i.e. assuming $z\approx 6.2897$. The very good quality of the X-Shooter spectrum allowed us also to measure the column density of a very weak Fe $\scriptstyle\rm II$ 2382 Å̊$\mathring{\rm A}$ absorption line and to determine an upper limit to the transition of Al $\scriptstyle\rm II$ 1670 Å̊$\mathring{\rm A}$. The absorption lines and the Voigt profile fit for the three studied systems are shown in Fig. 5, 6 and 7; while the measured column densities are reported in Table 2.
@@
Appendix B Environments enriched by a PISN and another Pop III SN
Cosmological simulations predict that the first stars formed preferentially isolated, but they could also appear in small groups (see e.g. Skinner & Wise, 2020). Furthermore, in this work we aim at providing chemical diagnostics to study the fingerprints of the first stars in distant absorbers, including Lyman Limit systems and sub-DLAs. The nature of these systems is still unclear and they might be associated to gaseous filaments enriched by different low-mass galaxies hosting Pop III stars. For these reasons, in this Section we investigate how the chemical diagnostics for PISN-enriched environments are affected when considering the contribution of a second Pop III SN, i.e. further exploring the case $f_{\rm PopIII}=1$.
The results of our investigation are shown in Figs. 8-9, where we compare the chemical abundance ratios predicted for environments enriched by one PISN and another Pop III SN (i.e. two Pop III SNe and $f_{\rm PopIII}=1$) with those previously shown in Figs. 2-3. Clearly, these abundance ratios are widespread in our chemical diagnostics, covering the areas predicted for environments enriched by single PISNe, other Pop III SNe, and a mixture of PISNe and Pop II SN, i.e. the so-called confusion regions. Thus, absorbers in the confusions regions might be enriched by Pop III only ($f_{\rm PopIII}=1$) through a combination of a PISN and another Pop III SN, or predominantly enriched by a populations of normal Pop II SNe ($f_{\rm PopIII}\leq 0.5$). To uncover the nature of these absorbers we should use all the proposed diagnostics and, eventually, use a $\chi^{2}$ fitting procedure (see Appendix C) exploiting their overall abundance pattern.
It is important to note that in all our diagnostics the absorber J1018+0548 (black cross) is consistent for being enriched by a PISN and another Pop III SN, i.e., uniquely imprinted by Pop III stars. To unveil the properties of their progenitors, we need to compare its observed chemical abundance pattern with our theoretical predictions as we will explain in the next Section.
Figure 8: Same as Fig. 2, but the chemical abundances are inherited by: only PISNe (solid red), only other Pop III SNe (solid blue), a single PISN but also another Pop III SN in different proportions ($\geq 60\%$ PISN: solid orange, $\leq 50\%$ PISN: solid green). We show also the chemical abundances predicted for environments enriched mainly by PISNe but also by Pop II SNe ($f_{\rm PopIII}\geq 0.6$, dashed red) and mainly by Pop II SNe but also by PISNe ($f_{\rm PopIII}\leq 0.5$, dotted grey, see Figs. 2-3).
Figure 9: The same as Fig. 3, but with the additional models described in the caption of Fig. 8.
Appendix C Fitting procedure
$\chi^{2}=\sum_{i=1}^{N}\dfrac{(O_{i}-T_{i})^{2}}{\sigma^{2}_{O_{i}}+\sigma^{2}_{T_{i}}}+\sum_{i=N+1}^{N+U}\dfrac{(O_{i}-T_{i})^{2}}{\sigma^{2}_{T_{i}}}\times\Theta(O_{i}-T_{i})+\sum_{i=N+U+1}^{N+U+L}\dfrac{(O_{i}-T_{i})^{2}}{\sigma^{2}_{T_{i}}}\times\Theta(T_{i}-O_{i}),$ (C1)
where $O_{i}$ and $T_{i}$ are the observed and theoretically predicted abundance ratios [Xi/Fe] and $\sigma_{O_{i},T_{i}}$ are their respective errors. In order to account for the uncertainties on the stellar yields, we assume $\sigma_{T_{i}}=0.50$, as suggested by Nomoto et al. (2013) and Hartwig et al. (2018). $\Theta(x)$ is the Heaviside function. The reduced chi-square, $\chi^{2}_{\nu}$, is computed as $\chi^{2}/(N+L+U-M)$, where N is the number of data points with finite measurements, L (U) is the number of data points with lower (upper) limits and M is the number of free parameters of the model (M$=4$ with fixed mixing parameter of Pop III SNe). Finally, we define the relative chi square as $\chi^{2}_{rel}\coloneqq\chi^{2}_{\nu}/min(\chi^{2}_{\nu})$.
To find the best model for the enrichment of the absorption system at $z=3.39$ in the sightline of the QSO J1018+0548 (Saccardi et al., 2023), we look for the minimum $\chi^{2}_{\nu}$ by considering all the models presented in Sec. 2 (see Fig. 1, 2, 3), i.e. by accounting for an enrichment: (i) uniquely driven by a single PISN or a Pop III SN with different mass/energy; (ii) mainly ($\geq 60\%$ metals) driven by a single PISN/other Pop III SN with the contribution of Pop II SNe; (iii) only partially ($\leq 50\%$ metals) driven by a single PISN/other Pop III SN with the contribution of Pop II SNe. Furthermore, we also consider an imprint from two Pop III SNe with various masses and energies, including PISNe, with $f_{\rm PopIII}=1$ (i.e. excluding the contribution of Pop II SNe). As shown in Fig. 4 these models provide the best fits for the abundance pattern of J1018+0548.
$\chi^{2}_{\nu}$ $f_{\rm PopIII}$ $M_{1}\,\rm[M_{\odot}]$ $E_{1}\,\rm[10^{51}\,erg]$ $M_{2}\,\rm[M_{\odot}]$ $E_{2}\,\rm[10^{51}\,erg]$ $f\rm_{1}$
0.19 1.0 100.0 1.8 11.8 1.8 0.8
0.37 1.0 253.0 97.3 27.5 1.2 0.1
Table 3: Parameters of the best-fitting models of J1018+0548 shown in Fig. 4. $M_{1,2}$ and $E_{1,2}$ are the initial masses and SN explosion energies of the two Pop III stars polluting the environment. We indicate with $f_{1}$ the fraction of metals provided by the Pop III SN that explodes first. Thus, the fraction of metals contributed by the second SN can be computed as $f_{2}=1-f_{1}$.
Figure 10: Minimum relative reduced chi square computed fitting the abundances of J1018+0548 for models with different $f_{\rm PopIII}$ and different initial masses of the SN that explodes first, $M_{\rm PopIII}$ or $M_{1}$.
In Table 3, we present the parameters of the models displayed in Fig. 4, which correspond to an enrichment driven by: (i) two Pop III SNe with same explosion energy, $E_{\rm 1,2}=1.8\times 10^{51}$ erg, but very different masses, $M_{1}=100\rm M_{\odot}$ and $M_{2}=11.8\rm M_{\odot}$ ($\chi^{2}_{\nu}=0.19$); (ii) a very massive PISN of $M_{1}=253\rm M_{\odot}$ and a second Pop III SN with lower mass, $M_{2}=27.5\rm M_{\odot}$, and lower explosion energy, $E_{\rm 2}=1.2\times 10^{51}$ erg ($\chi^{2}_{\nu}=0.37$). The parameter $f_{1}$ in the Table represents the fraction of metals provided by the Pop III SN/PISN that explodes first, while the fraction contributed by the second one will be $f_{2}=1-f_{1}$.
In Fig. 10 we show the minimum $\chi^{2}_{rel}$ obtained by studying all models with different $f_{\rm PopIII}$ and mass of the SN that explodes first, $M_{\rm PopIII}$, regardless of all the other parameters, i.e. $E_{1},M_{2},E_{2},f_{1},M_{\rm PopII}$. The chi square shown for models with $f_{\rm PopIII}=1$ can be either for environments enriched by single Pop III SNe with $M=M_{\rm PopIII}$ or multiple Pop III SNe, where the first SN that explodes has $M_{1}=M_{\rm PopIII}$.
We can clearly see in the Figure that all models with $f_{\rm PopIII}<1$ have $\chi^{2}_{rel}>3$, apart from a specific model with $f_{\rm PopIII}=0.8$ (providing $\chi^{2}_{rel}=2.21$, $\chi^{2}=0.45$) that describes the combined pollution by a single Pop III SN with $M=\rm 100M_{\odot}$ and a normal core-collapse Pop II SN with $M=\rm 13M_{\odot}$. Thus, we can safely exclude that the absorber J1018+0548 has been predominantly imprinted by Pop II SNe. As shown in the Figure, the models with the lowest $\chi^{2}_{rel}$ are the ones with $f_{\rm PopIII}=1$, i.e., those uniquely polluted by Pop III SNe: both PISNe+Pop III SNe and two Pop III SNe. We chose to show in Fig. 4 the abundance patterns of the best models of these two categories, i.e. models that have $\chi^{2}_{rel}\leq 2$ and that account for an enrichment by two Pop III SNe or by one PISN and a Pop III SN.